OpenStudy (96mertcank96):

the measure of two angles are complementary are in the ratio 1:5 find the degree measure of each angle

OpenStudy (anonymous):

Complementary angles are two angles whose sum is 90 degrees. A ratio of 1: 5 can be expressed algebraically as 1x: 5x. So our equation is 1x+ 5x = 90 6x = 90 x = 15 So one angle is 1x = 1(15) = 15 degrees, and the second angle is 5x = 5(15) = 75 degrees.

OpenStudy (96mertcank96):

thank you!